# CoBRA: A Universal Strategyproof Confirmation Protocol for Quorum-based Proof-of-Stake Blockchains

### v2

@@ 21          $suffixindex++$
27
28procedure slash($validators$)
29    $blacklist\leftarrow blacklist\cup validators$
30
Rewards
Participation rewards are sourced from transaction fees and distributed to validators based on their contributions to block finalization. Rewards are proportional to the fraction of blocks with valid certificates proposed by each validator and are distibuted every $n$ rounds. Specifically, we define the interval of rounds $[kn+1,(k+1)n]$ for $k\in\mathbb{Z}$ as the leader window $W_{k}$. After the end of the window $W_{k}$ the total transaction fees from finalized transactions $F_{W_{k}}$, are allocated based on the proportion of unique proposers $P_{W_{k}}$ relative to $n$: $R_{W_{k}}=F_{{W_{k}}}\cdot\dfrac{P_{W_{k}}}{n}.$ To ensure that correct validators successfully propose blocks, rewards are distributed only if $P_{W_{k}}\geq n-f$, i.e., at least $n-f$ unique validators submit proposals within $n$ rounds. This guarantees censorship resilience, meaning that rational validators have no incentive to accept bribes from Byzantine validators to censor specific transactions or remain silent when a correct block proposer is present.
4.5 Extending Finalization with Strong Chains
Recall that validators finalize transaction volumes up to $D$ coins within a time window $2\Delta^{*}$. This is the lower bound in our threat model. When the system is not under attack (i.e., fewer faulty nodes), blocks often receive more than $12f+$ 1 signatures. We leverage the certificates with additional signatures to further limit forks and finalize larger transaction volumes without waiting for $2\Delta^{*}$.
To achieve this, we introduce the concept of $i$-strong chains [17]. A ledger is $i$-strong if it contains valid certificates signed by $2f+1+i$ unique validators within $2\Delta^{*}$. With at least $1+i$ correct validators, the maximum number of forks decreases to $f-i$, allowing secure finalization of higher transaction volumes. In particular, when a certificate has $12.5f+$ 1 signatures, we can define a strongest chain that can finalize infinite transaction volume (regardless of $D$).
Definition 9 ($i$-strong chain).
Consider a validator $p$ and its local data structure $\text{ledger}^{p}$ (Algorithm 1, line 1). Let $S$ represent the set of distinct signatures from validators included in the valid certificates of blocks within the set of blocks in $\text{ledger}^{p}$ that were output by the SMR protocol $\Pi$ within the last $2\Delta^{*}$. We say that $\text{ledger}^{p}$ is an $i$-strong chain, where $i=|S|-2f-1$.
Strong Chain finalization rule
Strongest Chain finalization rule
We additionally define the strongest chain, as an i-strong chain with $i>\frac{f+1}{2}$. More than half of the correct validators participate in the strongest chain. This guarantees that only one strongest chain can exist within $2\Delta^{*}$, and all correct validators will shift on it after detecting a fork, under this rule:
5 Analysis
@@
Combining both guarantees demonstrates that the composition of $\Pi$ and CoBRA forms a rationally robust protocol. To formalize this, we define the utility function of the rational validators.
@@
$\mathcal{M}(\mathcal{F},I)=\frac{\gamma\cdot C(I)}{R(\mathcal{F})},$
where $\gamma$ is the maximum number of forks that can be created by the coalition, $C(I)$ is an upper bound of coins finalized per fork during $I$ as specified by the protocol, and $R(\mathcal{F})$ returns the number of rational validators in the coalition [^7] To provide stronger guarantees, we assume that all the rewards from double-spending are distributed exclusively and equally to the rational validators..
@@
5.1 Lemmas for proving safety
In Lemma 1, we prove that once a correct validator finalizes a $2\Delta^{*}$-old block $b$, no conflicting block can exist. Thus, misbehaving validators can only double-spend in $2\Delta^{*}$-recent blocks.
@@ 5.1 Lemmas for proving safety
Case 1: $p_{j}$ triggers $b_{2}.\mathsf{finalize}$ at time $t<t^{*}$. Then $p_{j}$ must have received $QC(b_{2})$ at some time $t^{\prime}\leq t<t^{*}$. Since $p_{j}$ broadcasts $QC(b_{2})$, validator $p_{i}$ receives it by some time $localTime_{b_{2}}\leq t+\Delta^{*}<localTime_{b_{1}}+2\Delta^{*}$. At time $localTime_{b_{2}}$, $p_{i}$ detects the equivocation and slashes the misbehaving validators (line 13), resulting in a non-empty blacklist: $QC(b_{1})\cap QC(b_{2})\neq\emptyset$. Therefore, at time $localTime$, the finalization condition (line 19) is not satisfied, and $p_{i}$ will not finalize $b_{1}$ (line 20), contradicting the assumption.
Case 2: $p_{j}$ triggers $b_{2}.\mathsf{finalize}$ at time $t\geq t^{*}$. At time $t$ $p_{j}$ has received both $QC(b_{1})$ and $QC(b_{2})$. Let $t_{b_{2}}$ and $t_{b_{1}}$ be the times at which $p_{j}$ received $QC(b_{2})$ and $QC(b_{1})$, respectively, and assume without loss of generality that $t_{b_{1}}\geq t_{b_{2}}$. At time $t_{b_{1}}$, $p_{j}$ detects the equivocation and slashes the misbehaving validators (line 13), resulting in $QC(b_{1})\cap QC(b_{2})\neq\emptyset$. Therefore, at time $t$, the finalization condition (line 19) is not satisfied, and $p_{j}$ will not finalize $b_{2}$ (line 20), leading to a contradiction.
In Lemma 2, we show that when validators employ the stake-bounded finalization rule, any rational validator engaging in a forking attack can gain at most $D$ coins.
@@
  1. 1.
    Upper bound of double-spending: Any rational validator participating in the attack can double-spend at most an amount of coins equal to the stake $D$.
  2. 2.
@@
(1) Upper bound of double-spending. Misbehaving validators can double-spend only coins in conflicting blocks that are accepted by clients. Any inclusion proof $\pi_{tx}$ for a transaction $tx$ included in a block $b$ is accepted by some client $c$ only after witnessing at least $12f+$ 1 finality votes for $b$. Therefore at least one correct validator $p_{i}$ has finalized block $b$. Since correct validators do not trigger $finalize$ for conflicting blocks, and at least one correct validator is required to finalize a block, there can be at most $f+1$ conflicting ledgers. Moreover, according to Lemma 1, there can only be $2\Delta^{*}$-recent conflicting blocks.
Assume the worst case, where there are $f+1$ conflicting ledgers and let us fix any rational validator $p$ participating in $f+1-m$ conflicting ledgers $L_{i\in[f+1-m]}$ with $m\in[f+1]$. The coins that misbehaving validators double-spend on these conflicting ledgers is upper bounded by the total coins spent in conflicting blocks finalized by at least one correct validator. To bound this, we fix $L_{i}$ for some $i\in[f+1-m]$, and upper bound the spent coins in $2\Delta^{*}$-recent blocks that are finalized by at least one correct validator.
Let $Suffix^{t^{\prime}}_{L_{i}}$ be the set of all $2\Delta^{*}$-recent blocks with valid certificates in $L_{i}$ until time $t^{\prime}$. For every block $b\in Suffix^{t^{\prime}}_{L_{i}}$, $p_{i}$ updates the variable $suffixvalue$ with the total output sum of transactions in the block $b$ and only finalizes $b$ if $suffixvalue$ is at most $C=D$ (line 22). Therefore, the amount of coins spent in $L_{i}$ is at most $D$.
Now, we will upper bound the total amount of coins validator $p$ can double-spend in all the $f+1-m$ conflicting ledgers. First, at most $m+1$ correct validators participate in each ledger. We prove that by reaching a contradiction. Assume there is a ledger with at least $m+2$ correct validators. Then, in the remaining $f-m$ ledgers, there must be at least one correct validator. Therefore, there are, in total, at least $m+2+f-m=f+2$ correct validators in the system. We reached a contradiction. Since there are at most $f$ byzantine and $m+1$ correct validators per ledger, at least $f-m$ rational validators participate in each ledger. Therefore, at least $f-m$ rational validators can double-spend at most $C=D$ coins in $f-m$ ledgers, and since they distribute the benefit equally, each validator double-spends $max_{m\in[f+1]}\frac{f-m}{f-m}D=D$ coins.
@@
Now, we argue that any misbehaving validator will be detected and forfeit their stake.
@@
Assume that the validators in $\mathcal{F}$ fork the system at time $t$ [^8] We refer to the first round when distinct correct validators observe conflicting blocks with valid certificates. As before, we use the local clock of any validator to refer to time elapsed between events. resulting in $j\in[f+1]$ conflicting ledgers $L_{i\in\{1,...,j\}}$. Let us fix any validator $p\in\mathcal{F}$. We will show that $p$ cannot withdraw its stake in any of the conflicting ledgers.

### v3

@@ 40    $suffixindex++$
47
48procedure slash($(validators,state)$)
49  return $state\leftarrow\textsc{slashAll}(validators,state)$ $\triangleright$ penalize every validator in the set
50
51procedure recover
52  broadcast all (s1)(s\!-\!1)–phase certificates which form finality certificates
53  continue after $2\Delta^{*}$
54  $\mathcal{L}\leftarrow\{L_{1},...,L_{j}\}$ $\triangleright$ conflicting finalized ledgers stored in finalitystorefinalitystore
55  $\mathcal{L}^{*}\leftarrow$ $Sort(\mathcal{L})$ $\triangleright$ lexicographically sort ledgers by the hash of their last block
56  $recovered\_state\leftarrow oldFinalizedBlock$
57  for pospos from $oldFinalizedBlock.height+1$ to $|ledger|$ do
58   if $ledger[pos]\in L$ for some $L\in\mathcal{L}$ continue
59   recovered_state.apply(ledger[pos])recovered\_state.apply(ledger[pos])
60  for all ledger $L\in\mathcal{L}$ do
61   for all block $b\in L$ do
62    recovered_state.apply(brecovered\_state.apply(b)
63  $recovered\_state\leftarrow$ slash(blacklist,recovered_stateblacklist,\;recovered\_state)
64  broadcast signed state commitment to recovered_staterecovered\_state
5.3. New Genesis
After reaching to a common state, validators produce a verifiable commitment to the recovered state, allowing new participants to trust the correct state. In particular, validators sign a state commitment (e.g. Merkle commitment) for the recovered state. To ensure that only honest validators can produce this certificate, they also attach the finality votes of misbehaving validators on conflicting blocks as evidence (line 64). Once at least $f+1$ validators sign this state commitment, along with the proofs of misbehavior, the resulting certificate constitutes the new genesis block.
Since at least $f+1$ validators have provably misbehaved–and these must be Byzantine, as CoBRA ’s finalization and slashing rules make equivocation unprofitable for rational validators–the remaining population of $n^{\prime}\leq 2f$ contains an honest majority, and at least $2n^{\prime}/3$ validators in total are honest or rational. Consequently, the system can continue running CoBRA, ensuring that no further equivocations occur.
6. Security Guarantees
In this section, we provide sketch proofs and defer the full analysis to Appendix B.
Safety and client-safety
Rational validators may collude with Byzantine validators to create conflicting blocks within the maximum message delay $\Delta^{*}$. However, such attacks are unprofitable: to double-spend, misbehaving validators must present inclusion proofs for blocks with finality certificates, ensuring they are finalized by at least one correct validator. During a fork, at most $fD$ coins can be double-spent in total, and any single rational validator gains at most $D$ coins, for all finalization rules—stake-bounded, strong, or strongest chain—as shown in Lemmas B.2B.4, and B.6, respectively.
Next, we argue that any misbehaving validator will be detected and forfeit their stake. Every correct validator observes all equivocations for blocks with finality certificates within $2\Delta^{*}$ and blacklists misbehaving validators via our slashing mechanism. Hence, misbehaving validators cannot withdraw their stake (Lemma B.3).
Finally, we conclude that rational validators have no incentive to engage in forking attacks, as any potential gain does not exceed their staked amount, which is subject to slashing. This ensures the protocol’s safety (Theorem B.11). Moreover, since clients accept inclusion proofs only for blocks with at least $2f+1$ finality votes, client-safety is also guaranteed (Theorem B.12).
Liveness and client-liveness
Since transaction fees are equally distributed among block proposers, rational validators have no incentive to withhold proposals or votes for valid blocks. However, Byzantine validators may attempt to bribe them to silence correct proposers and censor transactions.
Byzantine validators claim rewards only within windows where at least $2f+1$ unique proposers exist, ensuring at least one is correct. Once their rewards are depleted, rational validators resume proposing and voting for valid blocks (Lemma B.10). Consequently, a correct validator will eventually propose a block, ensuring liveness (Theorem B.11). Additionally, to claim their rewards, rational validators commit finality votes, guaranteeing client-liveness (Theorem B.12).
Recovery and economic restitution.
After a fork, all correct validators observe the equivocation and, within $2\Delta^{*}$, witness all blocks carrying conflicting finality certificates sent to clients. By CoBRA ’s finalization rules, the total value of coins that can be double-spent across conflicting finalized blocks is bounded by $fD$. Validators provably identify at least $f+1$ misbehaving validators, whose combined stake $(f+1)D$ covers all double-spent transactions. They then deterministically execute the transactions from conflicting blocks in the same order to agree on a common state and reimburse affected clients using the slashed stake (Theorem B.14).
7. Case-Study
These high participation rates are not unique to Cosmos. On Ethereum, the average daily participation has also consistently remained above 99% (Beaconscan, 2024). Thus, we conclude that CoBRA could be seamlessly integrated into production systems like Cosmos or Ethereum, enhancing system resilience without impacting block confirmation times.
8. Limitations and Extensions
Dynamic validator settings
We outline how to extend our protocol to the permissionless setting, where validators may dynamically join or leave the system. Similar to prior work on dynamic membership in BFT systems (Duan and Zhang, 2022), we assume that every membership change preserves the resilience assumptions required for safety and liveness. In our case, at any time at least one third of the active validators are honest, and between any two consecutive membership changes a quorum of at least two thirds—consisting of honest and rational validators—remains stable.
On-chain join requests. Similar to withdrawals, we treat join requests as on-chain state events. Once a join request is committed on-chain, validators update their local view of the validator set consistently. This guarantees that all existing validators agree on the current validator set, and the corresponding quorum size.
Challenges. Validator joining the system at a later time must obtain sufficient information to participate safely. In particular, late joiners must: (i) learn the current validator set, (ii) learn the current state of the system, (iii) detect past safety violations, ensuring that they neither enable malicious validators to withdraw their stake nor unknowingly contribute to the growth of a fork.
Membership and State Tracking. Late-joining validators can obtain the full chain and (i) execute the blocks to obtain the system state, (ii) follow the history of membership change events starting from the genesis block to reconstruct the current validator set. Optimizations such as introducing periodic checkpoints for the validator set and state commitments are left as future work.
Past equivocations. To ensure that late joiners do not mistakenly assist misbehaving validators, they must also learn whether any equivocation occurred in the past. We use the following mechanism:
  • Late joiner. A validator wishing to join in round $r$ broadcasts a join request and waits until round $r+2\Delta^{*}$ before participating fully. This waiting period ensures that it can receive any messages revealing past safety violations.
  • Existing validator. Any existing validator $p$ that receives the join request at some round $r^{\prime}\in\{r,r{+}\Delta^{*}\}$ responds to the joiner with all safety violations it is aware of up to round $r^{\prime}$. After that moment, $p$ treats the joining validator as active and interacts with it according to the protocol.
By round $r^{\prime}+\Delta^{*}\leq r+2\Delta^{*}$, the joining validator will have learned all equivocation evidence known to validator $p$ for rounds $\leq r^{\prime}$, and will learn about future inconsistencies during normal protocol participation.
Assessing transaction value
To address this, clients can include an estimated transaction value in their submissions. If the receiver agrees with the estimate, they accept the stake-based finality rule; otherwise, they wait for the full finality threshold $2\Delta^{*}$ (or a strong chain). However, this mechanism introduces a vulnerability: a malicious client can delay the finality of other transactions by regularly issuing, or falsely declaring, high-value transactions. To mitigate this, we impose a maximum insured value per transaction, ensuring sufficient throughput.
Bribing mechanism
Protocol Safety / Liveness Recovery Consistency Violations Recovery parameters Client Reimbursement
Classic BFT (Abraham et al., 2018; Buchman et al., 2018; Castro and Liskov, 2002) $n/3$ Byzantine
Accountable + Recoverable Partial Synchrony (Gong et al., 2025; Ranchal-Pedrosa and Gramoli, 2024) $n/3$ Byzantine $5n/9$ Byzantine 1 unbounded
$2n/3$ Byzantine 2
Accountable + Recoverable Synchrony (Lewis-Pye et al., 2025) $n/3$ Byzantine $5n/9$ Byzantine 1 $(f^{*}\Delta^{*},2\Delta^{*})$
$2n/3$ Byzantine 2
CoBRA (ours.)* $n/3$ Byzantine +n/3+\,n/3 rational $5n/9$ Byzantine +n/9+\,n/9 rational 1 $4\Delta^{*}$, no rollback
  • Columns: (i) resilience for safety/liveness, (ii) resilience for recovery, (iii–iv) number of consistency violations and recovery parameters ( time to achieve recovery, rollback of transactions finalized by clients), (v) client reimbursement. We denote by $\Delta^{*}$ the synchronous bound and by $f^{*}$ the actual number of Byzantine nodes.
    * Applying the recovery procedure of (Lewis-Pye and Roughgarden, 2025) on top of CoBRA ’s finalization rule (Sec. 4) guarantees client reimbursement against $2/3$ Byzantine adversary, at the cost of up to two safety violations, each requiring $O(f^{*}\cdot\Delta^{*})$ rounds.
Table 1. Comparison of classical BFT, accountable/recoverable consensus, and CoBRA.
Probabilistic network model.
In line with prior work, we assume an adversary that controls both malicious validators and network delays. We leave for future work models in which control over validators and the network is decoupled (Danezis et al., 2025) or network partitions are probabilistic (Guo et al., 2019). In such settings, attacks would succeed only with some probability, requiring the protocol to ensure that the expected utility of any attack for rational validators is strictly less than the stake they forfeit.
External incentives.
External incentives or attacks that are not observable within the blockchain ecosystem are incorporated by Byzantine behavior. By modeling a rational–Byzantine adversary, we capture a system that is resilient to both external incentives (Byzantine behavior), and internal incentives (rational behavior).
9. Related Work
Rational-Byzantine model in quorum-based protocol.
BAR (Aiyer et al., 2005) studies SMR with altruistic (honest), rational, and Byzantine parties. However, BAR is designed for general-purpose state machine replication and does not address the economic incentive structures or adversarial dynamics specific to payment systems. Moreover, BAR assumes that rational participants act independently and do not collude. Plenty of works considering rational and Byzantine parties also do not address collusions between participants (Amoussou-Guenou et al., 2019; McMenamin et al., 2021; Lev-Ari et al., 2019).
Abraham et al. (Abraham et al., 2006, 2008) introduce a solution concept which allows strategic collusion between rational and Byzantine players in the context of secret sharing and multi-party computation. Within this model, TRAP (Ranchal-Pedrosa and Gramoli, 2022) provides a protocol for one-shot consensus in the partially synchronous setting for $n>\max\left(3/2k+3f,2(k+f)\right)$. However, we show that this approach does not extend to multi-shot SMR in partial synchrony, as the total value that can be double-spent cannot be bounded. Assuming a synchronous network, we instead present an SMR protocol that tolerates up to $f+k<2/3$, improving upon TRAP’s resilience bound of $f+k<1/2$.
Accountability.
Several works focus on accountability in distributed protocols, including specific constructions (Civit et al., 2021; Buterin and Griffith, 2017; Shamis et al., 2022), analyzing forensic properties of well-known protocols (Sheng et al., 2021), and generic transformations that make protocols accountable (Civit et al., 2023). While these approaches ensure that misbehaving validators can eventually be detected, accountability alone does not guarantee rational resilience. If rational validators profit from double-spending is higher than their stake, they will misbehave even if eventually their stake is slashed. Moreover, this line of work does not address recovery after safety violations or client reimbursement. Our work closes these gaps by (i) identifying when accountability suffices to deter rational attacks and (ii) providing a recovery mechanism that ensures full reimbursement to clients when the Byzantine threshold is exceeded.
Economic safety
BLR24 (Budish et al., 2024b) examines economic safety in PoS blockchains, focusing on when slashing mechanisms can reliably punish equivocation after it occurs while ensuring honest validators retain their stake. They prove that slashing is impossible when more than $2f$ out of $3f+1$ validators misbehave, and propose a slashing mechanism when at most $2f$ do. Although our bounds align, as we also provide resilience up to $2f$ Byzantine and rational nodes, our focus is fundamentally different. BLR24 (Budish et al., 2024b) does not account for liquid coin transfers during SMR execution or explicitly model profit generation beyond slashing deterrents.
Recovery
Lewis-Pye et al. (Lewis-Pye and Roughgarden, 2025) study the impact of Byzantine adversaries controlling more than one third of the validators in quorum-based SMR. First, they show that adversaries controlling up to $5/9$ of validators can trigger at most one consistency violation, while adversaries controlling up to $2/3$ can trigger at most two. Moreover, they prove that in partial synchrony, recovery protocols cannot guarantee a bounded rollback window, implying that an unbounded number of finalized transactions may need to be reverted. As a consequence, such protocols (Ranchal-Pedrosa and Gramoli, 2024; Gong et al., 2025) cannot, in general, provide full client reimbursement and are therefore orthogonal to our approach, which targets client-level economic security.
Lastly, assuming a synchronous bound $\Delta^{*}$, they design a generic recovery protocol in which each recovery instance completes in $O(f^{*}\cdot\Delta^{*})$ rounds and rolls back transactions finalized within the previous $2\Delta^{*}$ rounds. In contrast, CoBRA provides a recovery procedure that completes within $4\Delta^{*}$, incurs at most one consistency violation, and ensures full reimbursement to all clients against an adversary controlling $5/9$ Byzantine plus $1/9$ rational validators. Notably, we can integrate the recovery procedure of (Lewis-Pye and Roughgarden, 2025) into CoBRA to tolerate up to $2/3$ Byzantine validators; full reimbursement would be still guaranteed, but recovery (up to two instances) would take $f^{*}\Delta^{*}$ rounds, essentially trading $1/9$ rational resilience for an extra $1/9$ Byzantine resilience at the cost of efficiency.
Sridhar et al. (Sridhar et al., 2023b) describe a gadget that can be layered onto synchronous SMR protocols to ensure that clients do not accept conflicting ledgers, but only under the assumption that clients actively communicate with one another and wait $3\Delta$ before finalizing transactions. In contrast, our clients remain silent: they simply accept blocks that carry finality votes. Any block meeting our finalization criteria can be finalized immediately. Moreover, (Sridhar et al., 2023b) does not specify recovery after safety violations and instead assumes that an honest majority is restored through an external process.
Algorithmic techniques
Decoupling the confirmation rules of clients and validators originates from the flexible BFT paradigm (Malkhi et al., 2019; Xiang et al., 2021; Neu et al., 2024) which separates the role of proposing and voting transactions (done by validators) with committing transactions (done by clients). We implement this separation using finality certificates for transactions and also leveraging certificates with high participation in our strong block finality rules. Finally, many protocols assume synchrony in quorum-based protocols for stronger guarantees, such as detecting equivocations (Sridhar et al., 2023a; Lewis-Pye and Roughgarden, 2025; Abraham et al., 2020), but in line with (Lewis-Pye and Roughgarden, 2025) our focus is to enable a transformation of partially synchronous protocols which can function in their assumed model without reliance on a synchrony bound.
Impossibilities
Finally, our impossibility results in partial synchrony align with (Tas et al., 2023; Budish et al., 2024b), which show that more than $1/3$ Byzantine validators can cause safety violations and avoid slashing once the security threshold is exceeded. We extend this by showing that the impossibility holds for any combination of Byzantine and rational participants exceeding $1/3$. In this setting, rational validators cannot be disincentivized from joining the attack.
Appendix A Rational attacks
In this section, we illustrate the challenges of extending (n,f)(n,f)–resilient SMR protocols towards (n,k,f)(n,k,f)–resiliency for $k\geq 1$.
Safety attacks
The attack. Rational validators can collude with adversarial validators to create two conflicting blocks $b_{1}$ and $b_{2}$ with valid inclusion proofs, where they spend $\alpha$ coins to two distinct subsets of clients $C_{1}$ and $C_{2}$. For clients in $C_{1}$ (or $C_{2}$), the $\alpha$ coins have not been spent yet, and therefore, rational validators can spend them again. In that case, we say that rational validators double-spend $\alpha$ coins. We present this attack in Lemma A.1, showing that $(n,f)$-resilient blockchains are vulnerable to double-spending.
Lemma 0 (Double-Spending Attack).
Consider a system of $n=f+k+h$ validators, where $f$ are Byzantine, $k$ are rational, and $h$ are correct validators. Let $\Delta^{*}$ denote the maximum message delay (either known or not before execution). For any $q\leq f+k+\lfloor h/2\rfloor$ and $f+k>1$ and any q–commitable protocol $\Pi$ that is $(n,f)$-resilient
$\Pi$ is vulnerable to double-spending attacks.
Proof.
Assume the leader $L_{r}$ proposing the next block at some round $r$ (w.l.o.g.) is malicious. If conflicting blocks already exist, the claim holds. Otherwise, all correct validators initially agree on the same ledger $T$. A malicious leader partitions the correct validators into two subsets, $H_{1}$ and $H_{2}$, where $|H_{1}|=\lfloor h/2\rfloor$ and $|H_{2}|=\lceil h/2\rceil$. Before the bound $\Delta^{*}$ elapses, the leader sends a conflicting block $b_{1}$ to $H_{1}$ and $b_{2}$ to $H_{2}$. Rational validators sign both $b_{1}$ and $b_{2}$ because they can present inclusion proofs for $b_{1}$ and $b_{2}$ to disjoint client sets $C_{1}$ and $C_{2}$ in exchange for off-chain assets. Each conflicting block gathers $q$ signatures since $f+k+\lfloor h/2\rfloor\geq q$ and $f+k+\lceil h/2\rceil\geq q$. Therefore, there is a valid certificate for the conflicting blocks $b_{1}$ and $b_{2}$. ∎
Liveness Attacks
Censoring transactions. In traditional Proof-of-Stake blockchains, the majority of transaction fees are awarded to the validator that successfully proposes a block, incentivizing the inclusion of as many transactions as possible. However, if a proposer fails to produce a valid block, the pool of unprocessed transactions grows, potentially increasing the rewards available to subsequent proposers. This dynamic creates an incentive for rational validators to launch censorship attacks (Kogias et al., 2016) against competing proposers to maximize their own rewards.
Bribery attacks. Byzantine validators can use their share of the rewards to bribe rational validators and censor all correct validators. As an example, we consider a 2f+12f+1–commitable SMR protocol with $f$ Byzantine, $f$ rational, and $f+1$ correct validators. To successfully launch the attack, Byzantine validators only need to bribe a single rational validator $p$, by distributing all their rewards to $p$. In that scenario, all the rest block proposers gather at most $2f$ signatures for their proposals. Consequently, only the byzantine validators along with $p$ would successfully propose a block. As a result, the adversary would manage to censor any transaction held by correct validators.
Appendix B Analysis
@@
Combining both guarantees demonstrates that the composition of $\Pi$ and CoBRA forms a coalition compliant protocol. To formalize this, we define the utility function of the rational validators.
@@
$\mathcal{M}(\mathcal{F},I)=\frac{\gamma\cdot C(I)}{R(\mathcal{F})}$
, where $\gamma$ is the maximum number of forks that can be created by the coalition, $C(I)$ is an upper bound of coins finalized per fork during $I$ as specified by the protocol, and $R(\mathcal{F})$ returns the number of rational validators in the coalition [^10] To provide stronger guarantees, we assume that all the rewards from double-spending are distributed exclusively and equally to the rational validators..
@@
B.1. Lemmas for proving safety
In Lemma B.1, we prove that once a correct validator finalizes a $2\Delta^{*}$-old block $b$, no conflicting block exists. Thus, conflicting block can exist only for $2\Delta^{*}$-recent blocks. We will use this lemma to establish the recovery guarantees of CoBRA.
@@ B.1. Lemmas for proving safety
Case 1: $p_{j}$ triggers $b_{2}.\mathsf{finalize}$ at time $t<t^{*}$. Then $p_{j}$ must have received $QC(b_{2})$ at some time $t^{\prime}\leq t<t^{*}$. Since $p_{j}$ broadcasts $QC(b_{2})$, validator $p_{i}$ receives it by some time $localTime_{b_{2}}\leq t+\Delta^{*}<localTime_{b_{1}}+2\Delta^{*}$. At time $localTime_{b_{2}}$, $p_{i}$ detects the equivocation and slashes the misbehaving validators (line 63), resulting in a non-empty blacklist: $QC(b_{1})\cap QC(b_{2})\neq\emptyset$. Therefore, at time $localTime$, the finalization condition (line 37) is not satisfied, and $p_{i}$ will not finalize $b_{1}$ (line 38), contradicting the assumption.
Case 2: $p_{j}$ triggers $b_{2}.\mathsf{finalize}$ at time $t\geq t^{*}$. At time $t$ $p_{j}$ has received both $QC(b_{1})$ and $QC(b_{2})$. Let $t_{b_{2}}$ and $t_{b_{1}}$ be the times at which $p_{j}$ received $QC(b_{2})$ and $QC(b_{1})$, respectively, and assume without loss of generality that $t_{b_{1}}\geq t_{b_{2}}$. At time $t_{b_{1}}$, $p_{j}$ detects the equivocation and slashes the misbehaving validators (line 63), resulting in $QC(b_{1})\cap QC(b_{2})\neq\emptyset$. Therefore, at time $t$, the finalization condition (line 37) is not satisfied, and $p_{j}$ will not finalize $b_{2}$ (line 38), leading to a contradiction.
In Lemma B.2, we show that when validators employ the stake-bounded finalization rule, any rational validator engaging in a forking attack can gain at most $D$ coins.
@@
  1. (1)
    Upper bound of double-spending: The coalition double spends up to $fD$ coins and any rational validator participating in the attack can double-spend at most an amount of coins equal to the stake $D$.
  2. (2)
@@
(1) Upper bound of double-spending. Misbehaving validators can double-spend only coins in conflicting blocks that are accepted by clients. Any inclusion proof $\pi_{tx}$ for a transaction $tx$ included in a block $b$ is accepted by some client $c$ only after witnessing a finality certificate $\mathcal{F}(b)$ including at least $2f+1$ finality votes for $b$. Therefore at least one correct validator $p_{i}$ has finalized block $b$. Since correct validators do not trigger finalizefinalize for conflicting blocks, and at least one correct validator is required to finalize a block, there can be at most $f+1$ conflicting ledgers.
Consider the time tt [^11] To refer at a specific time, we can fix the local clock of any validators, as only the time elapsed matters. where the first conflicting blocks with valid certificates are formed. Moreover, assume the worst case, where there are $f+1$ conflicting ledgers and let us fix any rational validator $p$ participating in f+1mf+1-m conflicting ledgers $L_{i\in[f+1-m]}$ with $m\in[f+1]$. The coins that misbehaving validators double-spend on these conflicting ledgers is upper bounded by the total coins spent in conflicting blocks finalized by at least one correct validator.
To bound this, we fix $L_{j}$ to be the longest ledger (i.e., the ledger with the most finalized blocks) at time $t$; if all ledgers have the same length, we fix $L_{j}$ for any $j\in[f+1-m]$. At time $t+\Delta^{*}$ every honest validator will have witnessed the equivocation and stop voting for new blocks. To upper bound the double spent coins during $t,t+\Delta^{*}]$ we fix some $L_{i}$ for any $i\in[f+1-m],i\neq j$ and upper bound the coins that finalized by at least one correct validator in $L_{i}$.
Let $Suffix^{[t,t+\Delta^{*}]}_{L_{i}}$ be the set of all blocks with valid certificates in $L_{i}$ during $t,t+\Delta^{*}]$. For every block $b\in Suffix^{[t,t+\Delta^{*}]}_{L_{i}}$, $p_{i}$ updates the variable suffixvaluesuffixvalue with the total output sum of transactions in the block $b$ and only finalizes $b$ if suffixvaluesuffixvalue is at most $C=D$ (line 41). Therefore, the amount of coins spent in $L_{i}$ is at most $D$.
Now, we will upper bound the total amount of coins validator $p$ can double-spend in all the f+1mf+1-m conflicting ledgers. First, at most $m+1$ correct validators participate in each ledger. We prove that by reaching a contradiction. Assume there is a ledger with at least $m+2$ correct validators. Then, in the remaining fmf-m ledgers, there must be at least one correct validator. Therefore, there are, in total, at least m+2+fm=f+2m+2+f-m=f+2 correct validators in the system. We reached a contradiction. Since there are at most $f$ byzantine and $m+1$ correct validators per ledger, at least fmf-m rational validators participate in each ledger. Therefore, at least fmf-m rational validators can double-spend at most $C=D$ coins in fmf-m ledgers, and since they distribute the benefit equally, each validator double-spends $max_{m\in[f+1]}\frac{f-m}{f-m}D=D$ coins.
@@
Now, we argue that any misbehaving validator will be detected and forfeit their stake. We emphasize that we consider forks with conflicting finality certificates—not merely valid quorum certificates—as only these can be double-spent and yield a profit for rational participants.
@@
Assume that the validators in $\mathcal{F}$ fork the system at time tt [^12] We refer to the first round when distinct correct validators observe conflicting blocks with valid certificates. As before, we use the local clock of any validator to refer to time elapsed between events. resulting in $j\in[f+1]$ conflicting ledgers $L_{i\in\{1,...,j\}}$ with finality certificates. Let us fix any validator $p\in\mathcal{F}$. We will show that $p$ cannot withdraw its stake in any of the conflicting ledgers.